# A coupled-channel perspective analysis on bottom-strange molecular pentaquarks

### v1

@@ Abstract
At present work, we systematically study various bottom-strange molecular pentaquarks to search for possible bound states and resonances by adopting one-boson-exchange model within complex scaling method. According to our calculations, we predict several bound and resonant states for bottom baryon $Y_{b}(\Lambda_{b},\Sigma_{b})\bar{K}^{(*)}$ and $Y_{b}K^{(*)}$ systems. In particular, a bound state in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system may correspond to the particle $\Xi_{b}(6227)$. Meanwhile, the predicted bound state with $6303\sim 6269~{}\rm{MeV}$ in the $I(J^{P})=1/2(1/2^{-})\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system is flavor exotic and does not appear in the spectroscopy of conventional baryons, which provides a practical way to clarify the nature of particle $\Xi_{b}(6227)$. We highly hope that our proposals can offer helpful information for the future experimental searches.
@@
After the discovery of the $X(3872)$ on the Belle experiment Belle:2003nnu in 2003, a large amount of data has been accumulated in the past two decades in high energy collision experiments. In the mean while, a series of new phenomenology studies related to the $XYZ$ and $P_{c}/T_{cc}$ states have been reported Liu:2019zoy ; Liu:2015fea . A detialed investigation of those exotic hadron states provides new insights for decoding their internal structures, which may deepen our understanding of the nonperturbative properties of quantum chromodynamics (QCD). Since many new particles locate near the hadron-hadron thresholds, these states could be naturally interpreted as candidates of molecular states Chen:2022asf ; Dong:2017gaw ; Guo:2017jvc ; Karliner:2017qhf ; Zou:2021sha ; Mai:2022eur ; Zou:2013af . It is highly desirable to identify those molecules states out of lots of candidates and predict more ones for experimental searches, which will motivate the experimental search of such molecular states.
@@
In 2021, the LHCb collaboration reported two resonances, namely $X_{0}(2900)$ and $X_{1}(2900)$ in $D^{-}K^{+}$ invariant mass spectrum by analyzing the decay amplitude of the $B^{+}\to D^{+}D^{-}K^{+}$ decay channel LHCb:2020pxc ; LHCb:2020bls . Since these two states are located near the $\bar{D}^{*}K^{*}$ and $\bar{D}_{1}^{*}K^{*}$ threshold, they are regarded as hadronic molecules candidates Chen:2020aos ; He:2020btl ; Burns:2020epm ; Agaev:2020nrc ; Xiao:2020ltm ; Kong:2021ohg ; Ke:2022ocs . Recently, in the analysis of $D_{s}^{+}\pi^{+}$ and $D_{s}^{+}\pi^{-}$ invariant mass spectrum, the LHCb collaboration has observed two new peaks $T_{c\bar{s}0}^{0}(2900)$ and $T_{c\bar{s}0}^{++}(2900)$, whose masses and widths are $M(T_{c\bar{s}0}^{0})=2892\pm 14\pm 15$ $\rm MeV$, $\Gamma=119\pm 26\pm 12$ $\rm MeV$ and $M(T_{c\bar{s}0}^{++})=2921\pm 17\pm 19$ $\rm MeV$, $\Gamma=137\pm 32\pm 14$ $\rm MeV$ LHCb:2022sfr ; LHCb:2022lzp , respectively. Given their near-threshold behaviors and quantum numbers, these two $T_{c\bar{s}0}^{0(++)}$ states are proposed as isovector $D^{*}K^{*}$ molecules with $J^{P}=0^{+}$ Agaev:2022eyk ; Wang:2023hpp ; An:2022vtg .
@@
Until now, most molecular candidates were observed in the charm sector, while the experimental observations in the bottom sector are still scarce. In 2006, the DØ Collaboration announced a narrow structure, referred to as the $X(5568)$ in $B_{s}^{0}\pi^{\pm}$ channel D0:2016mwd . Then, the LHCb collaboration investigated the $B_{s}^{0}\pi^{\pm}$ invariant mass spectrum, but no significant signal is found LHCb:2016dxl . Later, the ATLAS, CDF, and CMS collaborations ATLAS:2018udc ; CDF:2017dwr ; CMS:2017hfy released similar results. Meanwhile, the $X(5568)$ has been theoretically discussed in previous works Chen:2016ypj ; Xiao:2016mho ; Agaev:2016urs , and can not be assigned as an isovector $BK$ molecular state. In 2021, the LHCb collaboration reported two states in $B^{\pm}K^{\pm}$ mass spectrum, which are named as $B_{sJ}(6063)$ and $B_{sJ}(6114)$. If the missing photo from the $B^{*\pm}\to B^{\pm}\gamma$ was taken into consideration, the masses and widths were measured to be $B_{sJ}(6109):M=6108.8\pm 1.1\pm 0.7\rm MeV\text{ and }\Gamma=22\pm 5\pm 4\rm MeV$ and $B_{sJ}(6158):M=6158\pm 4\pm 5\rm MeV\text{ and }\Gamma=72\pm 18\pm 25\rm MeV$ LHCb:2020pet , respectively. In theory, the $B_{sJ}(6158)$ was widely investigated in the literature Kong:2021ohg . Some of the existing works suggested that $B_{sJ}(6158)$ can be interpreted as a $\bar{B}K^{*}$ molecular state with $I(J^{PC})=0(1^{+})$. Also, several works showed that the existence of $\bar{B}^{(*)}K^{(*)}(B^{(*)}\bar{K}^{(*)})$ molecular states are allowed Kolomeitsev:2003ac ; Guo:2006fu ; Guo:2006rp ; Sun:2018zqs .
@@
It can be seen that numerous exotic hadronic molecular states containing heavy quarks have been observed experimentally. The heavy quark symmetry is supposed to have been proven to play a significant role in predicting undiscovered states and understanding their production mechanisms, which intrigues several theoretical studies of it Asanuma:2023atv ; Tanaka:2024siw ; Wang:2023eng ; Sakai:2023syt . In a previous work, the author investigated open charm molecular counterpart of the newly $T_{c\bar{s}}^{a0(++)}$ composed of $Y_{c}(\Lambda_{c},\Sigma_{c})$ and strange meson $K^{(*)}$ interactions by adopting one boson exchange model. From their estimations, there can exist some bound states corresponding to the new observation $T_{c\bar{s}}^{a0(++)}$ Chen:2022svh . According to the heavy quark symmetry, on the bottom sector, the light diquark in the heavy baryons $\Sigma_{b}/\Lambda_{b}$ has the same color structure as $\bar{q}$, as shown in Figure 1. If $B_{sJ}(6158)$ can be explained as a $\bar{B}K^{*}$ molecular state with $I(J^{PC})=0(1^{+})$, there should also exist possible isoscalar $\bar{B}^{(*)}K^{(*)}$ molecular state. Under the circumstances, it is natural to conjecture whether there could exist possible open bottom molecular pentaquarks. Moreover, it is worth mentioning that, in 2018, the LHCb reported a peak in both $\Lambda_{b}^{0}K^{-}$ and $\Xi_{b}^{0}\pi^{-}$ invariant mass spectra named $\Xi_{b}(6227)$ LHCb:2018vuc . However, until now, whether the $\Xi_{b}(6227)$ should be accommodated into traditional baryon $\lambda-$mode $P-$wave $\Xi_{b}^{\prime}$ with $J^{P}=3/2^{-}$ and $5/2^{-}$ Chen:2018orb ; He:2021xrh ; Wang:2018fjm ; Cui:2019dzj or molecular state pure $\Sigma_{b}\bar{K}$ with $J^{P}=1/2^{-}$ is still on discussion Huang:2018bed ; Zhu:2020lza ; Mutuk:2024elj . Thus, it is urgent and necessary to explore the possibility of $\Xi_{b}(6227)$ being a molecular states and predict more bottom-strange molecular pentaquark candidates for future experiments.
@@
Recently, we systematically study the hidden bottom molecular tetraquark with complex scaling method by adopting one-boson-exchange(OBE) model Song:2024ngu , at present work, utilizing the same formulism, we systematically study various bottom-strange molecular pentaquarks to search for possible bound states and resonances by adopting within complex scaling method Aguilar:1971ve ; Balslev:1971vb ; Moiseyev:1998gjp ; Ho:1983lwa ; He:2021xrh and Gaussian expansion method Hiyama:2003cu ; Hiyama:2018ivm . For bottom baryon $Y_{b}(\Lambda_{b},\Sigma_{b})$ and anti-strange meson $\bar{K}^{(*)}$ interactions, our calculations demonstrate that some bound and resonant states are reveled. For instance, in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system, we obtain a bound state below $\Sigma_{b}\bar{K}$ threshold that can be regarded as the particle $\Xi_{b}(6227)$. Meanwhile, we extend our study to $Y_{b}K^{(*)}$ systems, and find two flavor exotic bound states, which can be searched in future experiments.
[…]
@@ II.1 The effective interactions
In this work, we adopt the one-boson-exchange model to describe the interaction between the hadrons and analyze the formation mechanisms of molecular states. The chrial symmetric interacting Lagrangian which corresponds to the coupling between a bottom baryon and a light mesons, can be constructed as Liu:2011xc
@@
The axial current $A_{\mu}$, vector current $\mathcal{V}_{\mu}$, and the vector meson field strength tensor $F^{\mu\nu}(\rho)$ are defined by
@@
respectively. Here, $\xi=\text{exp}({P}/f_{\pi})$ and $\rho_{ba}^{\mu}=ig_{V}{V}_{ba}^{\mu}/\sqrt{2}$. The $\mathcal{B}_{\bar{3}}$, $\mathcal{B}_{\mu}=-\sqrt{\frac{1}{3}}(\gamma_{\mu}+\nu_{\mu})\gamma^{5}\mathcal{B}_{6}+\mathcal{B}_{6\mu}^{*}$, ${P}$, and ${V}$ denote the matrices of ground state of singly heavy baryons multiplets in $\bar{3}_{F}$, $6_{F}$, light pseudoscalar and vector mesons, respectively, whose explicit form read
$\mathcal{B}_{6}$ $=$ $\begin{pmatrix}\Sigma_{b}^{+}&\frac{\Sigma_{b}^{-}}{\sqrt{2}}\\ \frac{\Sigma_{b}^{-}}{\sqrt{2}}&\Sigma_{b}^{0}\end{pmatrix},\quad\mathcal{B}_{\bar{3}}=\begin{pmatrix}0&\Lambda_{b}^{0}\\ -\Lambda_{b}^{0}&0\end{pmatrix},$
$P$ $=$ $\begin{pmatrix}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&\pi^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}\end{pmatrix},\quad V=\begin{pmatrix}\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&\rho^{+}\\ \rho^{-}&-\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}\end{pmatrix}.$
@@
$V_{\Lambda_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}^{*}}$ $-\frac{1}{6}\frac{g_{4}g_{VVP}}{f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda_{0},m_{\pi 0},r)-\frac{1}{6\sqrt{2}}\frac{\lambda_{I}g_{V}g}{m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda_{0},m_{\rho 0},r)$
$V_{\Sigma_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}^{*}}$ $+\frac{1}{2}l_{S}g_{\sigma}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}Y(\Lambda,m_{\sigma},r)+\frac{g_{1}g_{VVP}}{6\sqrt{2}f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})\mathcal{G}(I)Y(\Lambda,m_{\pi},r)-\frac{g_{1}g_{VVP}}{18\sqrt{2}f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda,m_{\eta},r)$
$+\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\mathcal{G}(I)Y(\Lambda,m_{\rho},r)+\frac{\lambda_{S}g_{V}g}{8\sqrt{3}m_{\Sigma_{b}}}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\mathcal{G}(I)\nabla^{2}Y(\Lambda,m_{\rho},r)-\frac{\lambda_{S}g_{V}g}{24\sqrt{3}m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})\mathcal{G}(I)Y(\Lambda,m_{\rho},r)$
$-\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}Y(\Lambda,m_{\omega},r)-\frac{\lambda_{S}g_{V}g}{8\sqrt{3}m_{\Sigma_{b}}}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\nabla^{2}Y(\Lambda,m_{\omega},r)+\frac{\lambda_{S}g_{V}g}{24\sqrt{3}m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda,m_{\omega},r)$
@@
$-\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\omega},r)+\frac{1}{24m_{\Sigma_{b}}}\lambda_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\nabla^{2}Y(\Lambda,m_{\omega},r)$
@@
$V_{\Sigma_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}}$ $-\frac{g_{1}g\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})}{24\sqrt{2}f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{G}(I)Y(\Lambda_{2},m_{\pi 2},r)+\frac{g_{1}g}{72\sqrt{2}f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{2},m_{\eta 2},r)$
$+\frac{\lambda_{S}g_{V}g_{VVP}}{6\sqrt{3}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})\mathcal{G}(I)Y(\Lambda_{2},m_{\rho 2},r)-\frac{\lambda_{S}g_{V}g_{VVP}}{6\sqrt{3}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{2},m_{\omega 2},r)$
With the above Lagrangian at hand, one can obtain the relevant potentials straightforwardly by using the Breit approximation. The effective potential in momentum space reads

### v2

@@ Abstract
In the present work, we systematically study various bottom-strange molecular pentaquarks to search for possible bound states and resonances by adopting a one boson exchange model within the complex scaling method. According to our calculations, we predict several bound and resonant states for bottom baryon $Y_{b}(\Lambda_{b},\Sigma_{b})\bar{K}^{(*)}$ and $Y_{b}K^{(*)}$ systems. In particular, a bound state in the $I(J^{P})=1/2(1/2^{-})\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system may correspond to the particle $\Xi_{b}(6227)$. Meanwhile, the predicted bound state with mass in the range of $6303\sim 6269~\rm{MeV}$ in the $I(J^{P})=1/2(1/2^{-})\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system is flavor exotic and does not appear in the spectroscopy of conventional baryons, which may help clarify the nature of $\Xi_{b}(6227)$. We hope that our study can provide useful guidance for future experimental searches.
@@
After the discovery of the $X(3872)$ by the Belle experiment Belle:2003nnu in 2003, a large amount of data has been accumulated over the past two decades in high energy collision experiments. Meanwhile, a series of new phenomenological studies related to the $XYZ$ and $P_{c}/T_{cc}$ states have been reported Liu:2019zoy ; Liu:2015fea . A detailed investigation of those exotic hadron states provides new insights into decoding their internal structures, which may deepen our understanding of the nonperturbative properties of quantum chromodynamics (QCD). Since many new particles are located near the hadron-hadron thresholds, these states could be naturally interpreted as candidates for molecular states Chen:2022asf ; Dong:2017gaw ; Guo:2017jvc ; Karliner:2017qhf ; Zou:2021sha ; Mai:2022eur ; Zou:2013af . It is highly desirable to identify those molecular states out of the numerous candidates and predict more of them for experimental searches, which will motivate the experimental search for such molecular states.
@@
In 2021, the LHCb Collaboration reported two resonances, namely $X_{0}(2900)$ and $X_{1}(2900)$, in the $D^{-}K^{+}$ invariant mass spectrum by analyzing the decay amplitude of the $B^{+}\to D^{+}D^{-}K^{+}$ decay channel LHCb:2020pxc ; LHCb:2020bls . Since these two states are located near the $\bar{D}^{*}K^{*}$ and $\bar{D}_{1}^{*}K^{*}$ thresholds, they are regarded as hadronic molecule candidates Chen:2020aos ; He:2020btl ; Burns:2020epm ; Agaev:2020nrc ; Xiao:2020ltm ; Kong:2021ohg ; Ke:2022ocs . Recently, in the analysis of the $D_{s}^{+}\pi^{+}$ and $D_{s}^{+}\pi^{-}$ invariant mass spectra, the LHCb Collaboration has observed two new peaks, $T_{c\bar{s}0}^{0}(2900)$ and $T_{c\bar{s}0}^{++}(2900)$, whose masses and widths are $M(T_{c\bar{s}0}^{0})=2892\pm 14\pm 15$ $\rm MeV$, $\Gamma=119\pm 26\pm 12$ $\rm MeV$ and $M(T_{c\bar{s}0}^{++})=2921\pm 17\pm 19$ $\rm MeV$, $\Gamma=137\pm 32\pm 14$ $\rm MeV$ LHCb:2022sfr ; LHCb:2022lzp , respectively. Given their near-threshold behaviors and quantum numbers, these two $T_{c\bar{s}0}^{0(++)}$ states are proposed as isovector $D^{*}K^{*}$ molecules with $J^{P}=0^{+}$ Agaev:2022eyk ; Wang:2023hpp ; An:2022vtg .
@@
Until now, most molecular candidates were observed in the charm sector, while the experimental observations in the bottom sector are still scarce. In 2006, the DØ Collaboration announced a narrow structure, referred to as the $X(5568)$, in the $B_{s}^{0}\pi^{\pm}$ channel D0:2016mwd . Then, the LHCb Collaboration investigated the $B_{s}^{0}\pi^{\pm}$ invariant mass spectrum, but no significant signal was found LHCb:2016dxl . Later, the ATLAS, CDF, and CMS Collaborations ATLAS:2018udc ; CDF:2017dwr ; CMS:2017hfy released similar results. Meanwhile, the $X(5568)$ has been theoretically discussed in previous works Chen:2016ypj ; Xiao:2016mho ; Agaev:2016urs ; Guo:2016nhb ; Albaladejo:2016eps ; Albaladejo:2016lbb , and cannot be assigned as an isovector $BK$ molecular state. In 2021, the LHCb Collaboration reported two states in the $B^{\pm}K^{\pm}$ mass spectrum, which are referred to as $B_{sJ}(6063)$ and $B_{sJ}(6114)$. If the missing photon from the $B^{*\pm}\to B^{\pm}\gamma$ was taken into consideration, the masses and widths were measured to be $B_{sJ}(6109):M=6108.8\pm 1.1\pm 0.7~\rm MeV$ and $\Gamma=22\pm 5\pm 4~\rm MeV$, and $B_{sJ}(6158):M=6158\pm 4\pm 5~\rm MeV$ and $\Gamma=72\pm 18\pm 25~\rm MeV$ LHCb:2020pet , respectively. In theory, the $B_{sJ}(6158)$ has been widely investigated in the literature Kong:2021ohg . Some of the existing works suggested that $B_{sJ}(6158)$ can be interpreted as a $\bar{B}K^{*}$ molecular state with $I(J^{PC})=0(1^{+})$. Also, several works showed that the existence of $\bar{B}^{(*)}K^{(*)}(B^{(*)}\bar{K}^{(*)})$ molecular states is allowed Kolomeitsev:2003ac ; Guo:2006fu ; Guo:2006rp ; Sun:2018zqs .
@@
It can be seen that numerous exotic hadronic molecular states containing heavy quarks have been observed experimentally. The heavy quark symmetry is supposed to have been proven to play a significant role in predicting undiscovered states and understanding their production mechanisms, which has stimulated several theoretical studies Asanuma:2023atv ; Tanaka:2024siw ; Wang:2023eng ; Sakai:2023syt . In a previous work, the authors investigated the open-charm molecular counterparts of the newly observed $T_{c\bar{s}}^{a0(++)}$ composed of $Y_{c}(\Lambda_{c},\Sigma_{c})$ baryons and strange mesons $K^{(*)}$, by adopting the one-boson exchange model. From their estimations, some bound states may exist that correspond to the newly observed $T_{c\bar{s}}^{a0(++)}$ Chen:2022svh . According to heavy quark symmetry, in the bottom sector, the light diquark in heavy baryons $\Sigma_{b}/\Lambda_{b}$ has the same color structure as a light antiquark $\bar{q}$, as shown in Figure 1. If the $B_{sJ}(6158)$ can be explained as a $\bar{B}K^{*}$ molecular state with $I(J^{PC})=0(1^{+})$, one should also expect the existence of isoscalar $\bar{B}^{(*)}K^{(*)}$ molecular state. Under these circumstances, it is natural to conjecture whether open-bottom molecular pentaquarks could also exist. Moreover, it is worth mentioning that, in 2018, the LHCb Collaboration reported a peak in both the $\Lambda_{b}^{0}K^{-}$ and $\Xi_{b}^{0}\pi^{-}$ invariant mass spectra, named $\Xi_{b}(6227)$ LHCb:2018vuc . However, until now, whether the $\Xi_{b}(6227)$ should be accommodated into traditional baryon $\lambda$-mode $P$-wave $\Xi_{b}^{\prime}$ baryon with $J^{P}=3/2^{-}$ or $5/2^{-}$ Chen:2018orb ; He:2021xrh ; Wang:2018fjm ; Cui:2019dzj , or instead as a pure molecular state $\Sigma_{b}\bar{K}$ with $J^{P}=1/2^{-}$ Huang:2018bed ; Zhu:2020lza ; Mutuk:2024elj ; Ozdem:2021vry is still on discussion. Thus, it is both urgent and necessary to explore the possibility of $\Xi_{b}(6227)$ being a molecular state and to predict more bottom-strange molecular pentaquark candidates for future experimental searches.
@@
Recently, we systematically studied the hidden bottom molecular tetraquark with complex scaling method by adopting one-boson-exchange (OBE) model Song:2024ngu . In the present work, utilizing the same formalism, we systematically study various bottom-strange molecular pentaquarks to search for possible bound states and resonances by adopting the complex scaling method Aguilar:1971ve ; Balslev:1971vb ; Moiseyev:1998gjp ; Ho:1983lwa and Gaussian expansion method Hiyama:2003cu ; Hiyama:2018ivm . For bottom baryon $Y_{b}(\Lambda_{b},\Sigma_{b})$ and anti-strange meson $\bar{K}^{(*)}$ interactions, our calculations demonstrate that some bound and resonant states are revealed. For instance, in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system, we obtain a bound state below $\Sigma_{b}\bar{K}$ threshold that can be regarded as the particle $\Xi_{b}(6227)$. Meanwhile, we extend our study to $Y_{b}K^{(*)}$ systems, and find two flavor exotic bound states, which can be searched for in future experiments.
The rest of this paper is organized as follows. We briefly introduce the formalism of effective interactions and complex scaling method in Sec. II. In Sec. III, we present the numerical results and discussions for the $Y_{b}$$K^{(*)}$ and $Y_{b}$$\bar{K}^{(*)}$ systems. Finally, we summarize in Sec. 5.
@@ II.1 The effective interactions
In this work, we adopt the one-boson-exchange model to describe the interactions between the hadrons and analyze the formation mechanisms of molecular states. The chiral symmetric interacting Lagrangian, which corresponds to the coupling between a bottom baryon and a light meson, can be constructed as Liu:2011xc
@@
The axial current $A_{\mu}$, the vector current $\mathcal{V}_{\mu}$, and the vector meson field strength tensor $F^{\mu\nu}(\rho)$ are defined as follows
@@
respectively. Here, $\xi=\text{exp}({P}/f_{\pi})$ and $\rho_{ba}^{\mu}=ig_{V}{V}_{ba}^{\mu}/\sqrt{2}$. The $\mathcal{B}_{\bar{3}}$, $\mathcal{B}_{\mu}=-\sqrt{\frac{1}{3}}(\gamma_{\mu}+\nu_{\mu})\gamma^{5}\mathcal{B}_{6}+\mathcal{B}_{6\mu}^{*}$, P{P}, and V{V} denote the matrices of the ground states of singly heavy baryon multiplets in $\bar{3}_{F}$, $6_{F}$, light pseudoscalar, and vector mesons, respectively, whose explicit forms read
$\mathcal{B}_{6}$ $=$ $\begin{pmatrix}\Sigma_{b}^{+}&\frac{\Sigma_{b}^{0}}{\sqrt{2}}&\frac{\Xi_{b}^{(^{\prime},*)0}}{\sqrt{2}}\\ \frac{\Sigma_{b}^{0}}{\sqrt{2}}&\Sigma_{b}^{-}&\frac{\Xi_{b}^{(^{\prime},*)-}}{\sqrt{2}}\\ \frac{\Xi_{b}^{(^{\prime},*)0}}{\sqrt{2}}&\frac{\Xi_{b}^{(^{\prime},*)-}}{\sqrt{2}}&\Omega_{b}^{(*)-}\end{pmatrix},\quad\mathcal{B}_{\bar{3}}=\begin{pmatrix}0&\Lambda_{b}^{0}&\Xi_{b}^{0}\\ -\Lambda_{b}^{0}&0&\Xi_{b}^{-}\\ -\Xi_{b}^{0}&-\Xi_{b}^{-}&0\\ \end{pmatrix}$, (7)
$P$ $=$ $\begin{pmatrix}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&K^{0}\\ K^{-}&\bar{K}^{0}&\frac{2}{\sqrt{6}}\eta\\ \end{pmatrix}$$\quad$ (8)
$V=\begin{pmatrix}\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&\rho^{+}&K^{*+}\\ \rho^{-}&-\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&K^{*0}\\ K^{*-}&\bar{K}^{*0}&\phi\\ \end{pmatrix}$. (9)
@@
$V_{\Lambda_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}^{*}}$ $\frac{1}{6}\frac{g_{4}g_{VVP}}{f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda_{0},m_{\pi 0},r)-\frac{1}{6\sqrt{2}}\frac{\lambda_{I}g_{V}g}{m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda_{0},m_{\rho 0},r)$
$V_{\Sigma_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}^{*}}$ $\frac{1}{2}l_{S}g_{\sigma}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}Y(\Lambda,m_{\sigma},r)+\frac{g_{1}g_{VVP}}{6\sqrt{2}f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})\mathcal{G}(I)Y(\Lambda,m_{\pi},r)-\frac{g_{1}g_{VVP}}{18\sqrt{2}f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda,m_{\eta},r)$
$\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\mathcal{G}(I)Y(\Lambda,m_{\rho},r)+\frac{\lambda_{S}g_{V}g}{8\sqrt{3}m_{\Sigma_{b}}}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\mathcal{G}(I)\nabla^{2}Y(\Lambda,m_{\rho},r)-\frac{\lambda_{S}g_{V}g}{24\sqrt{3}m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})\mathcal{G}(I)Y(\Lambda,m_{\rho},r)$
$\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}Y(\Lambda,m_{\omega},r)-\frac{\lambda_{S}g_{V}g}{8\sqrt{3}m_{\Sigma_{b}}}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\nabla^{2}Y(\Lambda,m_{\omega},r)+\frac{\lambda_{S}g_{V}g}{24\sqrt{3}m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda,m_{\omega},r)$
@@
$\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\omega},r)+\frac{1}{24m_{\Sigma_{b}}}\lambda_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\nabla^{2}Y(\Lambda,m_{\omega},r)$
@@
$V_{\Sigma_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}}$ $\frac{g_{1}g\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})}{24\sqrt{2}f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{G}(I)Y(\Lambda_{2},m_{\pi 2},r)+\frac{g_{1}g}{72\sqrt{2}f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{2},m_{\eta 2},r)$
$\frac{\lambda_{S}g_{V}g_{VVP}}{6\sqrt{3}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})\mathcal{G}(I)Y(\Lambda_{2},m_{\rho 2},r)-\frac{\lambda_{S}g_{V}g_{VVP}}{6\sqrt{3}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{2},m_{\omega 2},r)$
Before applying the potentials in coordinate space, we adopt the Breit approximation, a standard approach within the one-boson-exchange (OBE) framework. In this approximation, the relativistic scattering amplitude is reduced to an instantaneous effective potential in momentum space, which is valid in the low-energy regime with small momentum transfer. This treatment enables us to extract spin-dependent interaction terms, such as central, spin-spin, tensor, and spin-orbit components. For a more detailed discussion, we refer the reader to Refs. Breit:1929zz ; Breit:1930zza . Thus, the effective potential in momentum space reads